# Predictive AI with External Knowledge Infusion for Stocks

### v1

@@ 4.2 Sequential Embedding Layer
Considering the strong temporal dynamics of stock markets, it is intuitive to regard the historical status of a stock as the most influential factor in predicting its future trend. Therefore, we first apply a sequential embedding layer to capture the sequential dependencies in the historical price data and technical indicators. Due to the significant advancements and advantages demonstrated by Transformers in processing sequential data, we choose to utilize a Transformer model to learn the sequential embeddings. Consequently, we feed the historical time series data $\mathcal{P}_{i}^{T}$ of stock $i$ at time step $T$ into the Transformer model, which generates the sequential embedding $s_{i}^{T}$ for each stock as $s_{i}^{T}=\mathsf{Transformer}\left(\mathcal{P}_{i}^{T}\right).$ The Transformer model processes the entire historical sequence simultaneously, capturing long-range dependencies effectively. We employed the encoder-only architecture of the Transformer (Vaswani et al., 2017) to acquire sequential embeddings.
@@ 4.3 Temporal Process Embeddings
We then consider relations on the dynamic knowledge graph as events of a heterogeneous Hawkes process on the graph and learn node and relation embeddings using the Heterogeneous Hawkes Process for Dynamic Heterogeneous Graph Embedding (HPGE) algorithm (Ji et al., 2021). For event $(h,r,t,\tau)$, we consider the intensity function $\tilde{\lambda}(h,r,t,\tau)$ defined as
$\mu_{r}(h,t)+\gamma_{1}\sum_{r^{\prime},t^{\prime}\in\mathcal{N}_{r^{\prime}}(h)}m_{1}(h,r,t,\tau,r^{\prime},t^{\prime},\tau^{\prime})$
$\qquad\qquad+\gamma_{2}\sum_{r^{\prime},t^{\prime}\in\mathcal{N}_{r^{\prime}}(t)}m_{2}(h,r,t,\tau,r^{\prime},t^{\prime},\tau^{\prime}),$
where $\mu_{r}$ is the relation type $r$-specific base intensity function between $h$ and $t$, and $m_{1}$ and $m_{2}$ are the mutual intensities between the entities, summing over all relation types $r^{\prime}$ and past $r^{\prime}$-neighborhoods.
@@
where $\mathbf{W}_{e}$ is a dense layer and the activation function $\sigma$ is LeakyReLU. The embeddings and weights are optimized by maximizing the intensity function for positive samples and minimizing the intensity function for negative samples. The loss function and the mutual information terms are as defined in Ji et al. (2021).
The obtained node embeddings from the sequential layer and the temporal process model are concatenated to create new node embeddings $E_{s,\text{tpp}}=\left(s_{i}\|e_{\text{tpp},i}\right)_{i}$.
@@ 4.4 Relational Embeddings
Figure 5: [Best viewed in color]The experimental setup for model training illustrated across 24 phases. In each phase, the model undergoes a training period of 250 days, followed by a validation period of 50 days and, subsequently, a testing period of 100 days.
To obtain the relational embedding, each entity $i$ in the knowledge graph is assigned a learnable embedding $e^{n}_{i}$ randomly initialized from $\mathcal{N}(0,1)$. The node embeddings $E_{s,\text{tpp}}$ are added to the embeddings of nodes corresponding to the selected stock assets. Together, this creates an entity feature matrix $\mathbf{e}^{n}=\left(e^{n}_{i}\right)_{i=1,\dots,N}$.
Similarly, features for the set of relation types are defined as embeddings $\left(e^{r}_{j}\right)_{j=1,\dots,M}$, where $M$ is the number of relation types, and are initialized from $\mathcal{N}(0,1)$. Each of these is concatenated with the corresponding relation embedding $e^{r}_{\text{tpp},j}$ obtained from the temporal process embedding method and projected to its original dimension to obtain relational embeddings $\mathbf{e}^{r}=\mathbf{W}_{r,\text{tpp}}\cdot\left[e_{r}\|e^{r}_{\text{tpp}}\right]$.
Additionally, embeddings are defined for the month $e^{\tau}_{m}$, day $e^{\tau}_{d}$ and hour $e^{\tau}_{h}$, with weights initialized from $\mathcal{N}(0,1)$. These are then summed together to obtain the temporal embedding for the start timestamp $\tau_{s}$, denoted by $\mathbf{e}^{\tau}$.
These node embeddings $\mathbf{e}^{n}$, along with edge attributes $\mathbf{e}^{e}=\mathbf{e}^{r}+\mathbf{e}^{\tau}$ are passed through two layers of Heterogeneous Edge-featured Graph Attention (HEAT) (Mo et al., 2021) and two linear layers to obtain the final combined relational embedding $E_{h}$ for all stocks.
@@ 4.5 Prediction Layer
The obtained embeddings are passed through a prediction layer to obtain a vector of ranking score
@@ 4.6 Loss Function
The loss function is defined as the sum of four terms as follows.
@@
$\mathcal{L}_{2}$ is the listwise loss function ApproxNDCG (Bruch et al., 2019) comparing the actual and predicted ranks, $\mathcal{L}_{3}$ is a binary cross-entropy loss between the prediction and an indicator variable whether the stock price moved in the upward direction, and finally, $\mathcal{L}_{4}$ is another cross-entropy loss between the predicted ranking score and an indicator whether the stock is in the top 5 with respect to the IRR metric. The detailed expressions for each of the above losses are provided in Appendix A.
@@ 5 Experimental Setup
Table 3: Comparison of the obtained metrics of TA-HKGE with the baseline models. The datasets NASDAQ100, S&P500, and NIFTY500 are represented as N100, S&P, and NIFTY. The best result for each dataset and holding period is highlighted in bold. It can be noted that TA-HKGE performed better in most metrics.
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 TRANSF N100 0.088 25.027 0.874 0.214 1.014 0.08 22.498 0.09 0.315 5.556
S&P 0.062 16.868 0.748 0.143 0.181 0.065 17.9 0.03 0.208 1.103
NIFTY 0.108 33.477 0.484 0.152 0.389 0.098 28.173 -1.02 0.222 1.372
STHGCN N100 0.054 14.654 0.761 0.209 0.861 0.082 22.971 0.117 0.309 4.931
S&P 0.054 14.841 0.667 0.139 0.194 0.052 13.969 -0.02 0.204 0.8
NIFTY 0.055 14.982 0.62 0.21 0.82 0.069 19.114 0.06 0.31 4.74
STHAN N100 0.055 14.982 0.617 0.209 0.819 0.069 19.114 0.057 0.308 4.744
S&P 0.056 15.258 0.589 0.14 0.153 0.059 16.013 0.007 0.205 0.914
NIFTY 0.087 24.658 -1.32 0.15 0.347 0.112 32.582 -0.72 0.22 1.433
GCNKG N100 0.09 27.302 0.88 0.213 1.056 0.1 28.741 0.135 0.312 5.136
S&P 0.10 27.84 0.92 0.142 0.292 0.07 19.24 0.02 0.207 1.072
NIFTY 0.128 38.14 -0.621 0.148 0.278 0.109 31.538 -1.08 0.218 1.233
TA-HKGE N100 0.09 25.44 0.18 0.291 5.00 0.11 31.92 0.1 0.397 13.15
S&P 0.05 13.42 0.376 0.163 1.58 0.08 22.32 0.059 0.228 2.30
NIFTY 0.2 64.44 2.32 0.18 0.25 0.17 53.42 0.05 0.246 4.63
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
5 TRANSF N100 0.403 22.524 1.787 0.214 1.125 0.422 23.675 0.551 0.316 5.736
S&P 0.247 13.263 1.307 0.139 0.042 0.276 14.882 0.42 0.206 0.942
NIFTY 0.457 25.848 1.36 0.147 0.25 0.501 28.633 0.071 0.22 1.317
STHGCN N100 0.455 25.744 1.926 0.215 1.306 0.484 27.559 0.623 0.318 6.114
S&P 0.323 17.64 1.542 0.142 0.306 0.328 17.961 0.482 0.208 1.242
NIFTY 0.498 28.956 1.149 0.149 0.472 0.502 28.737 0.049 0.22 1.411
STHAN N100 0.466 26.507 2.09 0.214 1.236 0.489 27.917 0.611 0.318 6.197
S&P 0.45 25.6 2.08 0.144 0.5 0.363 20.028 0.517 0.209 1.217
NIFTY 0.564 33.529 -1.34 0.145 0.292 0.489 27.867 0.031 0.217 1.322
GCNKG N100 0.414 23.165 2.015 0.211 0.833 0.417 23.333 0.545 0.316 5.744
S&P 0.395 22.119 2.029 0.14 0.153 0.308 16.801 0.48 0.206 1.025
NIFTY 0.59 35.72 1.44 0.148 0.347 0.516 29.757 0.065 0.219 1.356
TA-HKGE N100 0.74 45.0 0.853 0.296 6.58 0.62 36.54 0.597 0.397 13.33
S&P 0.37 20.45 1.15 0.166 1.5 0.18 9.48 0.20 0.222 2.9
NIFTY 0.8 49.42 2.29 0.153 0.37 0.65 38.61 0.652 0.227 2.13
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
20 TRANSF N100 1.641 22.782 3.78 0.21 1.264 1.711 23.829 1.382 0.315 5.95
S&P 1.029 13.769 2.892 0.139 0.153 1.126 15.154 1.142 0.205 1.239
NIFTY 2.152 30.789 3.369 0.147 0.264 2.31 33.351 1.225 0.221 1.442
STHGCN N100 1.702 23.719 4.37 0.207 1.069 1.718 23.935 1.371 0.314 6.2
S&P 1.685 23.617 3.964 0.143 0.653 1.49 20.46 1.34 0.209 1.619
NIFTY 2.633 38.794 4.092 0.15 0.92 2.358 34.141 1.28 0.221 1.586
STHAN N100 1.742 24.328 4.338 0.208 0.931 2.013 28.548 1.54 0.316 6.819
S&P 1.79 25.13 4.2 0.144 0.847 1.376 18.846 1.165 0.211 1.567
NIFTY 2.95 44.49 4.52 0.147 0.486 2.388 34.657 1.33 0.22 1.592
GCNKG N100 1.504 21.345 3.848 0.211 1.417 1.404 19.654 1.126 0.319 6.372
S&P 1.172 16.001 3.196 0.138 0.181 1.168 15.794 1.085 0.206 1.253
NIFTY 2.632 38.978 4.109 0.147 0.569 2.47 36.06 1.294 0.22 1.575
TA-HKGE N100 3.51 54.46 6.067 0.289 4.87 2.0 28.34 0.718 0.391 12.99
S&P 1.5 20.63 4.4 0.166 1.08 1.56 21.53 0.907 0.235 3.53
NIFTY 6.65 125.06 6.585 0.150 1.42 1.55 21.38 1.88 0.236 2.67
@@
Three datasets are created containing a subset of stocks from NASDAQ and NIFTY, with datasets NASDAQ100 and S&P500 from the NASDAQ market and NIFTY500 from NSE, using Yahoo Finance. Each dataset consists of historical price movements and the constructed knowledge graph valid for each trading day. The historical price movements include features Close, Open, Low and High prices, and trading volume for each trading day from 30 January 2012 to 20 October 2022. The statistics are given in Table 2.
@@
The model is given as input a context window of stock data $X_{T}$ at time $T$, which consists of the historical stock price $\mathcal{P}^{\tau\in(T-W,T]}$ and the positive subgraph $G_{T}$ from the knowledge graph at the time point $T$. To avoid data leakage and minimize distribution drift, the values in the historical price window are normalized by dividing them by the last value in the previous window.
@@
For learning embeddings using the heterogeneous Hawkes process, we transformed the data tuples into the format (head, head_type, tail, tail_type, relation, timestamp), with a separate tuple generated for each month within the range of validity of the relation.
Table 4: TPP Shows the performance of the TA-HKGE model with a 750-day training period (750TR) and a 1000-day training period (1000TR).
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 750TR N100 0.112 32.47 0.652 0.268 2.5 0.121 35.67 0.428 0.375 12.12
S&P 0.004 0.92 0.179 0.154 0.25 0.066 18.02 0.306 0.225 2.62
NIFTY 0.139 41.83 -499.2 0.163 0.57 0.094 26.7 0.470 0.233 1.97
S&P 0.158 48.78 0.463 0.162 1.16 0.078 21.81 0.329 0.233 2.15
NIFTY 0.087 24.38 0.515 0.155 0.17 0.111 32.4 0.485 0.227 1.83
5 750TR N100 0.38 21.58 1.043 0.275 2.49 0.617 36.34 1.035 0.384 12.46
S&P 0.718 43.42 1.923 0.166 1.05 0.457 25.85 0.856 0.236 3.39
NIFTY 0.447 25.2 1.008 0.159 0.5 0.332 18.21 0.665 0.231 1.93
1000TR N100 0.411 22.97 0.984 0.279 3.27 0.645 38.25 1.006 0.390 13.81
S&P 0.142 7.4 0.597 0.169 1.03 0.621 36.65 1.433 0.244 4.14
NIFTY 0.901 57.1 2.102 0.157 0.71 0.638 37.76 1.181 0.23 2.16
20 750TR N100 0.378 4.87 1.536 0.258 2.17 1.456 19.98 1.698 0.3723 11.58
S&P 0.588 7.66 1.254 0.187 1.35 1.256 17.03 1.420 0.258 4.34
NIFTY 2.637 38.8 3.404 0.162 0.1 2.332 33.7 2.107 0.235 2.51
1000TR N100 1.334 18.18 1.433 0.265 4.01 2.082 29.64 2.440 0.373 11.35
S&P 0.907 12.05 1.767 0.162 0.11 1.853 26.02 1.945 0.237 3.6
NIFTY 0.613 8.0 0.565 0.147 0 1.679 23.34 1.724 0.219 1.33

### v2

@@ 5.1 Sequential Embedding Layer
Given the strong temporal dynamics of stock markets, it is intuitive to regard a stock’s historical performance as the most influential factor in predicting its future trend. Therefore, we first apply a sequential embedding layer to capture the sequential dependencies in the historical price data and technical indicators. Owing to the significant advancements and advantages demonstrated by transformers in processing sequential data, we chose to utilize a transformer model to learn the sequential embeddings. Consequently, we feed the historical time series data $\mathcal{P}_{i}^{T}$ of stock $i$ at time step $T$ into the Transformer model, which generates the sequential embedding $s_{i}^{T}$ for each stock as $s_{i}^{T}=\mathsf{Transformer}\left(\mathcal{P}_{i}^{T}\right).$ The Transformer model processes the entire historical sequence simultaneously, effectively capturing long-range dependencies. We employed the encoder-only architecture of the Transformer (Vaswani et al., 2017) to obtain sequential embeddings.
@@ 5.2 Temporal Process Embeddings
We then consider relations on the dynamic knowledge graph as events of a heterogeneous Hawkes process on the graph and learn node and relation embeddings using the Heterogeneous Hawkes Process for Dynamic Heterogeneous Graph Embedding (HPGE) algorithm (Ji et al., 2021). For event $(h,r,t,\tau)$, we consider the intensity function defined as
$\tilde{\lambda}(h,r,t,\tau)=\mu_{r}(h,t)+\gamma_{1}\sum_{r^{\prime},t^{\prime},\tau^{\prime}\in\mathcal{N}_{<\tau}(h)}m_{1}(h,r^{\prime},t^{\prime},\tau^{\prime})$
$\qquad\qquad+\gamma_{2}\sum_{r^{\prime},h^{\prime},\tau^{\prime}\in\mathcal{N}_{<\tau}(t)}m_{2}(t,r^{\prime},h^{\prime},\tau^{\prime}),$
where $\mu_{r}$ is the relation type $r$-specific base intensity function between $h$ and $t$, and $m_{1}$ and $m_{2}$ are the mutual intensities between the entities, summing over all relation types $r^{\prime}$ in the past neighborhoods of the nodes $h$ and $t$ with respective neighbors $t^{\prime}$ and $h^{\prime}$.
@@
where $\mathbf{W}_{e}$ is a dense layer, and the activation function $\sigma$ is LeakyReLU. The embeddings and weights are optimized by maximizing the intensity function for the positive samples and minimizing the intensity function for the negative samples. The loss function and mutual information terms are defined for the HPGE algorithm (Ji et al., 2021). The node embeddings obtained from the sequential layer and temporal process model are concatenated to create new node embeddings $E_{s,\text{tpp}}=\left(s_{i}\|e_{\text{tpp},i}\right)_{i}$.
@@ 5.3 Relational Embeddings
To obtain the relational embedding, each entity $i$ in the knowledge graph is assigned a learnable embedding $e^{n}_{i}$ randomly initialized from $\mathcal{N}(0,1)$. The node embeddings $E_{s,\text{tpp}}$ are added to the embeddings of the nodes corresponding to the selected stock assets. Together, this creates an entity feature matrix $\mathbf{e}^{n}=\left(e^{n}_{i}\right)_{i=1,\dots,N}$. Similarly, features for the set of relation types are defined as embeddings $\left(e^{r}_{j}\right)_{j=1,\dots,M}$, where $M$ is the number of relation types, and are initialized from $\mathcal{N}(0,1)$. Each of these is concatenated with the corresponding relation embedding $e^{r}_{\text{tpp},j}$ obtained from the temporal process embedding method and projected to its original dimension to obtain the relational embeddings
$\mathbf{e}^{r}=\mathbf{W}_{r,\text{tpp}}\cdot\left[e_{r}\|e^{r}_{\text{tpp}}\right].$
Additionally, embeddings are defined for the month $e^{\tau}_{m}$, day $e^{\tau}_{d}$ and hour $e^{\tau}_{h}$, with weights initialized from $\mathcal{N}(0,1)$. These are then summed to obtain the temporal embedding for the start timestamp $\tau_{s}$, denoted by $\mathbf{e}^{\tau}$. These node embeddings $\mathbf{e}^{n}$, along with edge attributes $\mathbf{e}^{e}=\mathbf{e}^{r}+\mathbf{e}^{\tau}$ are passed through two layers of Heterogeneous Edge-featured Graph Attention (HEAT) (Mo et al., 2021) and two linear layers to obtain the final combined relational embedding $E_{h}$ for all stocks.
@@ 5.4 Prediction Layer
The obtained embeddings are passed through a prediction layer to obtain a vector of ranking scores
@@
Table 2: Top 5 Performance comparison of TA-HKGE against baseline models, reported as averages over all phases of the corresponding datasets NASDAQ100, S&P500, and NIFTY500. Best results for each holding period are shown in bold.
$\Delta$ Model NASDAQ100 S&P100 NIFTY500
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 TRANSF 0.080 22.498 0.090 0.315 5.556 0.065 17.900 0.030 0.208 1.103 0.098 28.173 -1.020 0.222 1.372
STHGCN 0.082 22.971 0.117 0.309 4.931 0.052 13.969 -0.020 0.204 0.800 0.069 19.114 0.060 0.310 4.740
STHAN ‘ 0.069 19.114 0.057 0.308 4.744 0.059 16.013 0.007 0.205 0.914 0.112 32.582 -0.720 0.220 1.433
GCNKG 0.100 28.741 0.135 0.312 5.136 0.070 19.240 0.020 0.207 1.072 0.109 31.538 -1.080 0.218 1.233
CI-STHPAN 0.105 30.120 0.118 0.355 9.420 0.075 20.410 0.038 0.218 1.680 0.140 42.600 -0.420 0.232 2.980
TA-HKGE 0.110 31.920 0.100 0.397 13.150 0.080 22.320 0.059 0.228 2.300 0.170 53.420 0.050 0.246 4.630
5 TRANSF 0.422 23.675 0.551 0.316 5.736 0.276 14.882 0.420 0.206 0.942 0.501 28.633 0.071 0.220 1.317
STHGCN 0.484 27.559 0.623 0.318 6.114 0.328 17.961 0.482 0.208 1.242 0.502 28.737 0.049 0.220 1.411
STHAN 0.489 27.917 0.611 0.318 6.197 0.363 20.028 0.517 0.209 1.217 0.489 27.867 0.031 0.217 1.322
GCNKG 0.417 23.333 0.545 0.316 5.744 0.308 16.801 0.480 0.206 1.025 0.516 29.757 0.065 0.219 1.356
CI-STHPAN 0.520 30.180 0.570 0.355 9.240 0.240 12.900 0.330 0.214 1.860 0.585 34.120 0.320 0.223 1.740
TA-HKGE 0.620 36.540 0.597 0.397 13.330 0.180 9.480 0.200 0.222 2.900 0.650 38.610 0.652 0.227 2.130
20 TRANSF 1.711 23.829 1.382 0.315 5.950 1.126 15.154 1.142 0.205 1.239 2.310 33.351 1.225 0.221 1.442
STHGCN 1.718 23.935 1.371 0.314 6.200 1.490 20.460 1.340 0.209 1.619 2.358 34.141 1.280 0.221 1.586
STHAN 2.013 28.340 1.540 0.316 6.819 1.376 18.846 1.165 0.211 1.567 2.388 34.657 1.330 0.220 1.592
GCNKG 1.404 19.654 1.126 0.319 6.372 1.168 15.794 1.085 0.206 1.253 2.470 36.06 1.294 0.220 1.575
CI-STHPAN 1.720 24.180 0.920 0.355 9.640 1.360 18.740 0.980 0.220 2.420 2.020 28.900 1.560 0.228 2.050
TA-HKGE 2.000 28.548 0.718 0.391 12.990 1.560 21.530 0.907 0.235 3.530 1.550 21.380 1.880 0.236 2.670
@@ 5.5 Loss Function
The loss function is defined as the sum of the following four terms:
@@
where $\mathcal{L}_{2}$ is the listwise loss function ApproxNDCG (Bruch et al., 2019) comparing the actual and predicted ranks, $\mathcal{L}_{3}$ is a binary cross-entropy loss between the prediction and an indicator variable whether the stock price moved in the upward direction, and finally, $\mathcal{L}_{4}$ is another cross-entropy loss between the predicted ranking score and an indicator whether the stock is in the top 5 with respect to the IRR metric. The detailed expressions for each of the above losses are provided in Appendix A.
@@ 6 Experimental setup
6.1 Datasets
@@
Three datasets are created, containing a subset of stocks from NASDAQ and NIFTY, with datasets NASDAQ100 and S&P500 from the NASDAQ market and NIFTY500 from NSE, using Yahoo Finance. Each dataset consists of historical price movements and a constructed knowledge graph valid for each trading day. The historical price movements include features Close, Open, Low and High prices, and trading volume for each trading day from 30 January 2012 to 20 October 2022. The statistics are presented in Table 3.
@@
The model is given as input a context window of stock data $X_{T}$ at time $T$, which consists of the historical stock price $\mathcal{P}^{\tau\in(T-W,T]}$ and the positive subgraph $G_{T}$ from the knowledge graph at time point $T$. To avoid data leakage and minimize distribution drift, the values in the historical price window are normalized by dividing them by the last value in the previous window.
@@
To learn embeddings using the heterogeneous Hawkes process, we transformed the data tuples into the format (head, head_type, tail, tail_type, relation, timestamp), with a separate tuple generated for each month within the range of validity of the relation.